OpenStudy (anonymous):

integral from 0 to 2 of (5-((4-x^2)^(1/2))

hartnn (hartnn):

did you try to substitute \(x=2\sin u\) ?

hartnn (hartnn):

dx= ... ? 4-x^2 =... ?

OpenStudy (anonymous):

Yup. I got to the integral of 4cos^2x, but am a little stuck on this part...

hartnn (hartnn):

(5-2cos u)(2 cos u) du how did u just get 4 cos^2 u ?

hartnn (hartnn):

i get 10 cos u -4cos^2 u du

OpenStudy (anonymous):

hah, my bad, I pulled out the 5 as a separate integral

OpenStudy (anonymous):

but where you are is where I'm stuck

hartnn (hartnn):

cool so, cos^2 u write that in terms of cos2u, can u ?

hartnn (hartnn):

know the formulas for cos 2x ?

OpenStudy (anonymous):

I don't...there's a formula?

OpenStudy (anonymous):

I don't understand how I can write cos^2u in terms of cos(2u)

hartnn (hartnn):

yup \(\large \cos2x=\cos^2x-\sin^2x=2\cos^2x-1=1-2\sin^2x\) heard of them ?

hartnn (hartnn):

so, from cos2x = 2 cos^2x-1 could you isolate cos^2 x???

OpenStudy (anonymous):

You mean from the 4?

hartnn (hartnn):

i just mean, whether you can isolate cos^2x from cos2x = 2 cos^2x-1 ?

OpenStudy (anonymous):

I don't know, the concept is foreign to me,

hartnn (hartnn):

cos2x = 2 cos^2x-1 so, cos^2 x = (cos 2x+1)/2 just simple algebra

hartnn (hartnn):

\(\int 4\cos^2x dx = 4\int\dfrac{\cos2x+1}{2}dx \) and can you integrate cos 2x and 1 ?

OpenStudy (anonymous):

ahh okay I got it!

OpenStudy (anonymous):

I'm good from here, was just confused on that step!

OpenStudy (anonymous):

thanks !

hartnn (hartnn):

welcome ^_^ and \[ \begin{array}l\color{red}{\text{W}}\color{orange}{\text{E}}\color{#e6e600}{\text{L}}\color{green}{\text{C}}\color{blue}{\text{O}}\color{purple}{\text{M}}\color{purple}{\text{E}}\color{red}{\text{ }}\color{orange}{\text{t}}\color{#e6e600}{\text{o}}\color{green}{\text{ }}\color{blue}{\text{O}}\color{purple}{\text{p}}\color{purple}{\text{e}}\color{red}{\text{n}}\color{orange}{\text{S}}\color{#e6e600}{\text{t}}\color{green}{\text{u}}\color{blue}{\text{d}}\color{purple}{\text{y}}\color{purple}{\text{!}}\color{red}{\text{!}}\color{orange}{\text{}}\end{array} \]